10 1 2 3 4 5 6 7 8 9 10 1. What is the rise of the line? 2. What is the run of the line 3. Finally, what is the slope of the line?
Answer #1

1) Rise is the change in Y=3-2=1

2) Run is the change in X=4-2=2

3) slope is the rise over run=1/2